The mean time to assemble a piece of furniture is 52 minutes. Suppose the the assembly times are normally distributed with a standard deviation of 4.3 minutes. What percentage of assembly times fall between 47.7 minutes and 56.3 minutes?

Use the Empirical Rule

16%

68%

99.7%

95%
Answer #1

Let μ =52 and σ= 4.3 Here μ-σ=52-4.3 = 47.7 and μ + σ = 52 + 4.3=56.3 The required probability is p(47.7srs 56.3) ~ 68%, by e